Drive the dimension of
1. Surface tension
2. Coefficient of viscosity
3. Pressure
4. Universal gravitational force
5. Young modulus.
1
Expert's answer
2019-04-02T10:56:33-0400

1. Surface tension: since


"\gamma=\frac{1}{2}\frac{F}{L},"

"[\gamma]=\frac{M^1L^1T^{-2}}{L^1}=M^1T^{-2}."

2. Coefficient of viscosity:


"F=\mu A\frac{v}{y}, \space\space \implies\space\space \mu=\frac{yF}{vA},,"

"[\mu]=\frac{L^1\cdot M^1L^1T^{-2}}{L^1T^{-1}\cdot L^2}=M^1L^{-1}T^{-1}."

3. Pressure:


"p=F\/A,"

"[p]=\frac{M^1L^1T^{-2}}{L^2}=M^1L^{-1}T^{-2}."

4. Universal gravitational force:


"F=G\frac{M_1}{R^2}M_2=aM_2,"

"[F]=M^1L^1T^-2."

5. Young modulus:


"E=\sigma/\epsilon,"

"[E]=\frac{M^1L^{-1}T^{-2}}{1}=M^1L^{-1}T^{-2}."